15. Darwin has a capital of $ 7100.00 which he invests for 3 years at 7.6 % p. a. a)How much will he receive come maturity time if the interest is compounded annually? Round to nearest 100th. b)How much will he received come maturity time if the interest is compounded semi-annually? Round to nearest 100th . c)How much will he receive come maturity time if the interest is compounded quarterly? Round to nearest 100th .

*** no financial inputting in excel ** please show steps.
Answer #1

This question requires application of basic time value of money function, which is:

FV = PV * (1 + r)n

a) PV = $7100, n = 3, r = 7.6%

FV = 7100 * (1 + 7.6%)3

FV = 7100 * 1.2458

FV = $8,844.95

b) PV = $7100, n = 3 * 2 = 6 (semi-annual periods), r = 7.6%/2 = 3.8% (semi-annually)

FV = 7100 * (1 + 3.8%)6

FV = 7100 * 1.2508

FV = $8,880.60

c) PV = $7100, n = 3 * 4 = 12 (quarterly periods), r = 7.6%/4 = 1.9% (quarterly)

FV = 7100 * (1 + 1.9%)12

FV = 7100 * 1.2534

FV = $8,899.15